# Efficient Dataset Distillation via Diffusion-Driven Patch Selection for Improved Generalization

### v3

@@ 3.3 Visual Elements Aggregation
Although we have successfully selected the most representative patches, the overly complex data often introduce excessive randomness. To enhance the representative and reduce the randomness of the synthetic dataset, we perform an effective selecting strategy on the cropped patches as shown in the Stage II of Fig. 2. With diffusion model, we select DIFT [32] as the feature embedding for applying clustering, where noise at a specific time step is added to the latent of input image, and passed through a U-Net [24] with text condition to obtain the intermediate layer activations as features. On this basis, we use k-means [19] to cluster the selected patches. To better leverage clustering information, we have designed a ranking rule: 1) Intra-cluster. patches are ranked based on their distance to the centroid instead of corresponding $R(\mathbf{p}|c)$. 2) Inter-cluster. clusters are ranked based on the median $R(\mathbf{p}|c)$ of the patches within each cluster, the partial visualization is shown in Fig. 3. In practice, we fix the number of cluster centers at 32 and prioritize the selection of patches from the top 10 clusters under all compression settings to ensure both representativeness and diversity. A noteworthy point is that, unlike D4M, which dynamically updates cluster centers and generates samples accordingly, we utilize static clustering to obtain more class-relevant visual representations.
3.4 Fine-grained Image Reconstruction
When constructing the synthetic dataset for large-scale datasets (e.g., ImageNet-1K) using patches, we observed the following phenomenon: when the synthetic dataset size is limited with smaller IPC, the model’s ability to learn complex representations is constrained, thus the samples in synthetic dataset must represent the most prevalent patterns. By focusing on learning from these representative features, the model can capture essential discriminative information across classes. Therefore, when IPC $\leq 10$, we follow the experimental setup of RDED by concatenating multiple patches into a single image, allowing the model to focus on representative class features as shown in Fig. 4. However, as the IPC increases, complete feature is more likely to provide representations that closely align with statistical patterns, while excessively fragmented features tend to hinder network’s ability to improve accuracy and generalization, especially in deeper models with stronger learning capacity. Thus, we adopt the experimental settings of Minimax and D4M, where each patch is treated as an individual image. Experimental results demonstrate that our strategy yields effective outcomes.
Figure 3: Visualization of clustering. The top five patches from the top three clusters within the golden retriever class. Each row represents a single cluster.
3.5 Effective Label Calibration
Based on previous research [35, 31, 30] and preliminary experiments, we find that training large-scale synthetic datasets with hard labels can hinder the model from accurately learning useful information. In contrast, leveraging soft labels provided by the teacher model for the corresponding synthetic datasets can significantly improve the accuracy and generalization of the student model. Therefore, for a fair comparison, we also adopt FKD [27] which is utilized in RDED and D4M to align soft label during the training process, which can be formulated as:
$\mathbf{\theta}^{t+1}=\mathop{\arg\min}\limits_{\mathbf{\theta}\in\Theta}L_{KL}(T^{t}(\mathbf{x}),S_{\theta}^{t}(\mathbf{x})),$ (7)
where $T^{t}(x)$ and $S^{t}(x)$ represents the prediction for the distilled image $\mathbf{x}$ at training epoch $t$, $L_{KL}$ represents the Kullback–Leibler (KL) divergence.
Figure 4: Visualization comparison between RDED and ours under IPC=10. Our method can provide more complete class-specific representation.
@@ 3.5 Effective Label Calibration
Figure 5: Visualization comparison of images from the top ten classes of the ImageNet-1k dataset is presented (not cherry picked). From top to bottom, each row corresponds to the real dataset, RDED, Minimax, D4M, and ours under IPC=50. Compared to RDED, our method provides more comprehensive class-specific features. In contrast to the approach of directly using diffusion to generate synthetic datasets, our method fully adheres to the distribution of the real dataset and demonstrates a sharper focus on class-specific features.
ResNet-18 ResNet-101
Dataset IPC SRe2L Minimax D4M RDED OursC OursS SRe2L Minimax D4M RDED OursC OursS
10 29.4 $\pm$ 3.0 57.2 $\pm$ 0.8 57.4 $\pm$ 0.4 61.4 $\pm$ 0.4 62.5 $\pm$ 0.2 60.4 $\pm$ 0.3 23.4 $\pm$ 0.8 49.2 $\pm$ 0.3 49.6 $\pm$ 0.6 54.0 $\pm$ 0.4 56.4 $\pm$ 0.3 52.1 $\pm$ 0.4
ImageNette 50 40.9 $\pm$ 0.3 84.4 $\pm$ 0.5 84.8 $\pm$ 0.2 80.4 $\pm$ 0.4 83.4 $\pm$ 0.5 85.8 $\pm$ 0.4 36.5 $\pm$ 0.7 80.4 $\pm$ 0.9 80.6$\pm$ 0.4 75.0 $\pm$ 1.2 78.1 $\pm$ 0.4 82.4 $\pm$ 0.5
100 50.2 $\pm$ 0.4 90.1 $\pm$ 0.6 90.4 $\pm$ 0.7 89.6 $\pm$ 1.0 90.2 $\pm$ 0.2 91.6 $\pm$ 0.4 47.8 $\pm$ 0.6 88.7 $\pm$ 0.6 88.2 $\pm$ 1.4 88.2 $\pm$ 1.1 88.8 $\pm$ 0.2 90.4 $\pm$ 0.6
10 20.2 $\pm$ 0.2 38.1 $\pm$ 0.6 36.8 $\pm$ 1.2 38.5 $\pm$ 2.1 44.3 $\pm$ 0.5 40.1 $\pm$ 0.5 17.7 $\pm$ 0.9 34.6 $\pm$ 0.5 31.8 $\pm$ 1.2 31.3 $\pm$ 1.3 37.6 $\pm$ 0.4 34.8 $\pm$ 0.2
ImageWoof 50 23.3 $\pm$ 0.3 71.2 $\pm$ 0.6 71.4 $\pm$ 1.4 68.5 $\pm$ 0.7 70.1 $\pm$ 0.5 73.5 $\pm$ 1.4 21.2 $\pm$ 0.2 66.4 $\pm$ 0.6 67.2 $\pm$ 0.8 59.1 $\pm$ 0.7 65.2 $\pm$ 0.4 70.6 $\pm$ 0.2
100 37.6 $\pm$ 0.3 77.4 $\pm$ 0.2 78.5 $\pm$ 0.3 77.2 $\pm$ 0.6 76.9 $\pm$ 0.6 80.2 $\pm$ 0.8 34.4 $\pm$ 0.7 74.0 $\pm$ 0.4 73.4 $\pm$ 0.8 74.1 $\pm$ 0.5 73.8 $\pm$ 0.2 76.8 $\pm$ 0.4
10 9.5 $\pm$ 0.4 31.6 $\pm$ 0.7 33.4 $\pm$ 0.2 36.0 $\pm$ 0.3 37.0 $\pm$ 0.3 35.1 $\pm$ 0.3 6.4 $\pm$ 0.1 30.1 $\pm$ 0.7 33.1 $\pm$ 0.7 33.9 $\pm$ 0.1 36.5 $\pm$ 0.6 33.7 $\pm$ 0.4
ImageNet-100 50 27.0 $\pm$ 0.4 64.0 $\pm$ 0.5 62.7 $\pm$ 0.6 61.6 $\pm$ 0.1 64.2 $\pm$ 0.3 68.8 $\pm$ 0.6 25.7 $\pm$ 0.3 64.8 $\pm$ 0.6 65.3 $\pm$ 0.6 66.0 $\pm$ 0.6 67.3 $\pm$ 0.1 69.4 $\pm$ 0.6
100 39.4 $\pm$ 0.1 76.4 $\pm$ 0.3 76.8 $\pm$ 0.5 75.2 $\pm$ 0.3 75.2 $\pm$ 0.4 77.5 $\pm$ 0.2 38.2 $\pm$ 0.5 78.1 $\pm$ 0.7 77.6 $\pm$ 0.7 77.2 $\pm$ 0.3 76.9 $\pm$ 0.5 79.4 $\pm$ 0.6
10 21.3 $\pm$ 0.6 44.3 $\pm$ 0.3 41.9 $\pm$ 0.3 42.0 $\pm$ 0.1 43.9 $\pm$ 0.2 42.1 $\pm$ 0.3 30.9 $\pm$ 0.1 46.6 $\pm$ 0.7 44.6 $\pm$ 0.7 48.3 $\pm$ 1.0 51.2 $\pm$ 0.4 47.4 $\pm$ 0.3
ImageNet-1K 50 46.8 $\pm$ 0.2 56.6 $\pm$ 0.9 55.9 $\pm$ 0.3 56.5 $\pm$ 0.1 58.1 $\pm$ 0.3 59.4 $\pm$ 0.2 60.8 $\pm$ 0.5 62.6 $\pm$ 0.6 63.4 $\pm$ 1.0 61.2 $\pm$ 0.4 62.7 $\pm$ 0.2 65.4 $\pm$ 0.7
100 52.5 $\pm$ 0.5 58.6 $\pm$ 0.3 59.6 $\pm$ 0.4 60.5 $\pm$ 0.8 60.7 $\pm$ 0.3 61.8 $\pm$ 0.5 65.4 $\pm$ 0.2 67.3 $\pm$ 0.5 66.1 $\pm$ 0.9 65.4 $\pm$ 1.3 67.7 $\pm$ 0.1 70.0 $\pm$ 0.3
@@
Table 1: Performance comparison with methods applicable to large-scale datasets. Following the evaluation protocol of RDED, we use ResNet-18 as the teacher model for generating the dataset and evaluate on ResNet-18/101. OursC and OursS represent that each image in the synthetic datasets is concatenated patches and a single patch, respectively.
@@
Dataset IPC DSA IDM TESLA DATM D4M Ours SRe2L RDED Ours
CIFAR10 10 52.1 $\pm$ 0.5 58.6 $\pm$ 0.1 66.4 $\pm$ 0.8 66.8 $\pm$ 0.2 56.2 $\pm$ 0.4 48.8 $\pm$ 0.4 29.3 $\pm$ 0.5 37.1 $\pm$ 0.3 36.4 $\pm$ 0.2
50 60.6 $\pm$ 0.5 67.5 $\pm$ 0.1 72.6 $\pm$ 0.7 76.1 $\pm$ 0.3 72.8 $\pm$ 0.5 66.2 $\pm$ 0.3 45.0 $\pm$ 0.7 62.1 $\pm$ 0.1 61.0 $\pm$ 0.6
CIFAR100 10 32.3 $\pm$ 0.3 45.1 $\pm$ 0.1 41.7 $\pm$ 0.3 47.2 $\pm$ 0.4 45.0 $\pm$ 0.1 47.3 $\pm$ 0.4 27.0 $\pm$ 0.4 42.6 $\pm$ 0.2 41.5 $\pm$ 0.4
50 42.8 $\pm$ 0.4 50.0 $\pm$ 0.2 47.9 $\pm$ 0.3 55.0 $\pm$ 0.2 48.8 $\pm$ 0.3 57.6 $\pm$ 0.2 50.2 $\pm$ 0.4 62.6 $\pm$ 0.1 63.8 $\pm$ 0.3
TinyImageNet 10 - 21.9 $\pm$ 0.3 - 31.1 $\pm$ 0.3 35.4 $\pm$ 0.2 40.3 $\pm$ 0.2 16.1 $\pm$ 0.2 41.9 $\pm$ 0.2 40.2 $\pm$ 0.5
50 - 27.7 $\pm$ 0.3 - 39.7 $\pm$ 0.3 41.2 $\pm$ 0.5 49.2 $\pm$ 0.7 41.1 $\pm$ 0.4 58.2 $\pm$ 0.1 58.5 $\pm$ 0.2
@@ 4 Experiments
4.1 Experimental Settings
Datasets and Architectures.
To evaluate our proposed method on low-resolution datasets, we conduct experiments on CIFAR-10/100 [15] and TinyImageNet [16], with ConvNet [6] serving as the backbone. For high-resolution data, we perform evaluations on ImageNet-1K [4] and its widely used subsets, including ImageNet-100 [14], ImageNette and ImageWoof [1], utilizing ResNet-18 [9] as the backbone.
Baselines.
In addition to data-matching methods that are effective on low-resolution data: DSA [37], CAFE [33], and DATM [8], we consider recent methods applicable to large-scale datasets as strong competitors.
  • SRe2[35], by updating the synthetic dataset using the CE-loss of a pre-trained classifier and introducing soft label matching, was the first method to be feasibly applied to large-scale datasets.
  • RDED [31] significantly improved computational efficiency and performance by evaluating randomly cropped patches with a pre-trained classifier and concatenating them together to generate images.
  • Minimax [7] proposed fine-tuning DiT model pre-trained on ImageNet-1K by using the distribution of target dataset as a regularization to directly generate images.
  • D4[30] extended the range of diffusion models to pre-trained text-to-image Stable Diffusion, utilizing the cluster centers of latent to generate synthetic datasets.
Evaluation.
In line with previous work, we set IPC=10 and 50 for low-resolution datasets, and IPC=10, 50, and 100 for high-resolution datasets to enable more efficient learning. During the training process, we employ a fixed pre-trained classifier to provide soft labels. Subsequently, five randomly initialized networks are trained on the distilled dataset, with the average performance of these networks evaluated.
Implementation details.
When calculating $R(x|c)$ of images, we select a time-step range of [0.1, 0.7], whereas for the computation of DIFT, a time-step of 1.60 is employed. Across all IPC settings, we prioritize cropping the patches with resolution of 224x224 and selecting them from the top 10 ranked clusters after clustering. All the experiments were carried out using PyTorch on NVIDIA RTX-3090 GPUs.
@@ 4.2 Performance Improvements
For high-resolution datasets, we first resize the smallest edge of images to 256 pixels while maintaining height-width ratio. The resized images are then fed into the diffusion model to calculate $R(x|c)$ of the images and that of all patches with a 224x224 averaging pool kernel. As aforementioned, we adopt two settings for different compressing ratio to enhance the image construction performance. Specifically, for IPC $\leq$ 10, we select the top 4 patches from the top 10 clusters, resize each to 112x112, and concatenate them into a single image. When IPC $\geq$ 20, we directly use patches without resizing operation. For instance, when IPC = 100, we select 10 patches from the top ten clusters. If a cluster contains insufficient patches, we supplement the synthetic dataset by selecting the highest-scoring patches. As shown in Tab. 1, our proposed method demonstrates significant improvements across various datasets and IPC settings. In specific configurations, our approach exceeds baseline’s performance by more than 10%, further validating the superiority of our method.
Low-resolution results.
@@
For low-resolution datasets, we maintain the same setting as in RDED, utilizing the diffusion model to calculate the $R(x|c)$ of full-image without cropping, followed by clustering. We utilize ConvNet-3 and ConvNet-4 to generate soft labels for CIFAR-10/100 and TinyImageNet, respectively, and evaluate the synthetic datasets on the same architecture. As demonstrated in Tab. 2, the incorporation of the diffusion model yields performance on CIFAR-10 that is comparable to, or even exceeds, that of existing methods when applied within our proposed framework. On more complex datasets, such as CIFAR-100 and TinyImageNet, our approach outperforms all current optimization-based methods. Under high IPC settings, our method achieves SOTA performance, surpassing all prior techniques.
Efficient one-step distillation.

### v4

@@ 3.2 Visual Elements Aggregation
Although we have successfully selected the most representative patches, the overly complex data often introduce excessive randomness. To enhance the representativeness and reduce the randomness of the synthetic dataset, we perform an effective selecting strategy on the cropped patches as shown in the Stage II of Figure 2. With diffusion model, we select DIFT [tang2023emergent]for applying clustering, where noise at a specific time step is added to the latent of input image, and passed through a U-Net [ronneberger2015u] with text condition to obtain the intermediate layer activations as features. On this basis, we use K-means [lloyd1982least] to cluster the selected patches.
To better leverage clustering information, we have designed a ranking rule: 1) Intra-cluster. patches are ranked based on their distance to the centroid instead of corresponding $R(\mathbf{p}|c)$. 2) Inter-cluster. clusters are ranked based on the median $R(\mathbf{p}|c)$ of the patches within each cluster. In practice, we fix the number of cluster centers at 32 and prioritize the selection of patches from the top 10 clusters under all compression settings to ensure both representativeness and diversity. Furthermore, when IPC $\leq 10$, we follow the experimental setup of RDED by concatenating multiple patches into a single image. While for other settings, we treat each patch as an individual image.
@@
Figure 3: Visualization comparison of images from the top ten classes of the ImageNet-1k dataset is presented (not cherry picked). From top to bottom, each row corresponds to the real dataset, RDED, Minimax, D4M, and ours under IPC=50. Compared to existing methods, our method fully adheres to the distribution of the real dataset and demonstrates a sharper focus on class-specific features.
@@
Table 1: Performance comparison with methods applicable to large-scale datasets. Following the evaluation protocol of RDED, we use ResNet-18 as the teacher model for generating the dataset.
ResNet-18 DiT-ImageNet Stable Diffusion
Dataset IPC SRe2L RDED Minimax IGD D3HR D4M MGD3 Ours
10 29.4 $\pm$ 3.0 61.4 $\pm$ 0.4 57.2 $\pm$ 0.8 58.3 $\pm$ 0.2 58.1 $\pm$ 0.2 57.4 $\pm$ 0.4 58.9 $\pm$ 0.4 61.7 $\pm$ 0.4
ImageNette 50 40.9 $\pm$ 0.3 80.4 $\pm$ 0.4 84.4 $\pm$ 0.5 85.5 $\pm$ 0.3 85.7 $\pm$ 0.2 84.8 $\pm$ 0.2 85.2 $\pm$ 0.5 86.4 $\pm$ 0.2
100 50.2 $\pm$ 0.4 89.6 $\pm$ 1.0 90.1 $\pm$ 0.6 91.6 $\pm$ 0.5 90.9 $\pm$ 0.2 90.4 $\pm$ 0.7 91.0 $\pm$ 0.2 92.0 $\pm$ 0.3
10 20.2 $\pm$ 0.2 38.5 $\pm$ 2.1 38.1 $\pm$ 0.6 39.8 $\pm$ 0.1 39.3 $\pm$ 0.3 36.8 $\pm$ 1.2 38.2 $\pm$ 0.4 44.5 $\pm$ 0.4
ImageWoof 50 23.3 $\pm$ 0.3 68.5 $\pm$ 0.7 71.2 $\pm$ 0.6 72.8 $\pm$ 0.4 73.0 $\pm$ 0.2 71.4 $\pm$ 1.4 72.1 $\pm$ 0.5 73.5 $\pm$ 1.4
100 37.6 $\pm$ 0.3 77.2 $\pm$ 0.6 77.4 $\pm$ 0.2 78.3 $\pm$ 0.4 78.7 $\pm$ 0.3 78.5 $\pm$ 0.3 77.9 $\pm$ 0.2 79.6 $\pm$ 0.2
10 9.5 $\pm$ 0.4 36.0 $\pm$ 0.3 31.6 $\pm$ 0.7 33.2 $\pm$ 0.3 34.0 $\pm$ 0.2 33.4 $\pm$ 0.2 34.2 $\pm$ 0.5 36.8 $\pm$ 0.2
ImageNet-100 50 27.0 $\pm$ 0.4 61.6 $\pm$ 0.1 64.0 $\pm$ 0.5 64.2 $\pm$ 0.3 65.3 $\pm$ 0.2 62.7 $\pm$ 0.6 64.1 $\pm$ 0.3 67.5 $\pm$ 0.3
100 39.4 $\pm$ 0.1 75.2 $\pm$ 0.3 76.4 $\pm$ 0.3 77.8 $\pm$ 0.3 77.2 $\pm$ 0.2 76.8 $\pm$ 0.5 76.7 $\pm$ 0.3 78.6 $\pm$ 0.2
10 21.3 $\pm$ 0.6 42.0 $\pm$ 0.1 44.3 $\pm$ 0.3 45.5 $\pm$ 0.5 44.3 $\pm$ 0.3 27.9 $\pm$ 0.3 41.8 $\pm$ 0.2 46.1 $\pm$ 0.3
ImageNet-1K 50 46.8 $\pm$ 0.2 56.5 $\pm$ 0.1 58.6 $\pm$ 0.3 60.3 $\pm$ 0.4 59.4 $\pm$ 0.1 55.2 $\pm$ 0.3 57.2 $\pm$ 0.4 61.0 $\pm$ 0.3
100 52.5 $\pm$ 0.5 59.5 $\pm$ 0.8 60.2 $\pm$ 0.3 61.4 $\pm$ 0.3 62.5 $\pm$ 0.0 59.3 $\pm$ 0.4 61.7 $\pm$ 0.5 63.0 $\pm$ 0.2
@@
Dataset IPC DSA IDM TESLA DATM RDED Ours SRe2L D4M Ours
CIFAR10 10 52.1 $\pm$ 0.5 58.6 $\pm$ 0.1 66.4 $\pm$ 0.8 66.8 $\pm$ 0.2 50.2 $\pm$ 0.3 48.8 $\pm$ 0.4 29.3 $\pm$ 0.5 34.2 $\pm$ 0.3 36.4 $\pm$ 0.2
50 60.6 $\pm$ 0.5 67.5 $\pm$ 0.1 72.6 $\pm$ 0.7 76.1 $\pm$ 0.3 68.4 $\pm$ 0.1 66.2 $\pm$ 0.3 45.0 $\pm$ 0.7 60.9 $\pm$ 0.1 61.0 $\pm$ 0.6
CIFAR100 10 32.3 $\pm$ 0.3 45.1 $\pm$ 0.1 41.7 $\pm$ 0.3 47.2 $\pm$ 0.4 48.1 $\pm$ 0.3 47.3 $\pm$ 0.4 27.0 $\pm$ 0.4 40.8 $\pm$ 0.5 41.5 $\pm$ 0.4
50 42.8 $\pm$ 0.4 50.0 $\pm$ 0.2 47.9 $\pm$ 0.3 55.0 $\pm$ 0.2 57.0 $\pm$ 0.1 57.6 $\pm$ 0.2 50.2 $\pm$ 0.4 62.3 $\pm$ 0.4 63.8 $\pm$ 0.3
TinyImageNet 10 - 21.9 $\pm$ 0.3 - 31.1 $\pm$ 0.3 39.6 $\pm$ 0.1 40.3 $\pm$ 0.2 16.1 $\pm$ 0.2 35.1 $\pm$ 0.4 40.2 $\pm$ 0.5
50 - 27.7 $\pm$ 0.3 - 39.7 $\pm$ 0.3 47.6 $\pm$ 0.2 49.2 $\pm$ 0.7 41.1 $\pm$ 0.4 52.7 $\pm$ 0.4 58.5 $\pm$ 0.2
Teacher Network Student Network
ResNet-18 61.0 63.1 65.7 53.6 60.4 61.8 63.7
ResNet-50 52.2 64.8 66.7 48.9 56.2 51.3 60.1
ResNet-101 50.1 63.7 66.5 46.1 55.5 42.8 50.5
MobileNet-V2 55.2 62.5 65.0 55.3 60.5 57.2 62.3
EfficientNet-B0 45.8 59.1 61.1 45.0 58.4 51.7 60.3
Swin-T 34.4 51.2 52.7 35.1 44.1 51.1 50.2
ViT-B 31.2 46.9 49.3 32.7 40.0 41.3 46.2
@@ 4 Experiments
In this section, we demonstrate that our method outperforms existing approaches in both performance and computational efficiency. Deployment details and additional ablation studies are provided in the appendix.
@@ 4.1 Performance Improvements
High-resolution Results. For high-resolution datasets, we first resize the smallest edge of images to 256 pixels while maintaining height-width ratio. The resized images are then fed into the diffusion model to calculate $M(\mathbf{z}|c)$ of the images and that of all patches with a $224\times 224$ averaging pool kernel. For instance, when IPC = 100, we select 10 patches from the top ten clusters. If a cluster contains insufficient patches, we supplement the synthetic dataset by selecting the highest-scoring patches. As shown in Table 1, our proposed method demonstrates significant improvements across various datasets and IPC settings. In specific configurations, our approach exceeds baseline’s performance by more than 10%, further validating the superiority of our method.
@@
Low-resolution Results. For low-resolution datasets, we maintain the same setting as in RDED, utilizing the diffusion model to calculate the $M(\mathbf{z}|c)$ of full-image without cropping, followed by clustering, we then evaluate the synthetic datasets on the same architecture. As demonstrated in Table 2, the incorporation of the diffusion model yields performance on CIFAR-10 that is comparable to, or even exceeds, that of existing methods when applied within our proposed framework. On more complex datasets, such as CIFAR-100 and TinyImageNet, our approach outperforms all current optimization-based methods and achieves SOTA performance, surpassing all prior techniques.
4.2 Qualitative Analysis
To better understand the superiority of the datasets generated by our proposed method, we used t-SNE to visualize and analyze datasets produced by different methods. As shown in Figure 4, it is evident that the dataset generated by RDED [sun2024diversity] is overly concentrated around class centers because it originates from a pre-trained classifier. Similarly, the Minimax [gu2024efficient], which uses the DiT pre-trained on ImageNet-1K to generate images, also results in datasets that are overly concentrated in class centers. On the other hand, D4M [su2024d] generates images using SD, which is entirely unrelated to ImageNet-1K, resulting in datasets with excessive noise and outlier samples. In contrast to all the aforementioned methods, our approach achieves a good balance between diversity and representativeness by using SD for localization rather than generation. Furthermore, our method can still leverage diffusion models to generate images, which may lead to improved performance.
Figure 5: The separate time cost of individual tasks and total time cost of finishing all the tasks. Our method achieves a constant time cost due to one-step distillation process, which eliminates the need for repeated distillation.
4.3 Cross-architecture Generalization
Following previous work, we evaluate the performance of the synthetic dataset using different teacher-student combinations. It can be observed that, when using similar network architectures as teacher-student pairs (e.g., both CNNs), shallower networks generally yield better performance as teacher models, while the larger networks show stronger learning ability and obtain the best performance when the size of synthetic datasets increases.
We further compare the performance of the synthetic dataset on extremely different network architectures, e.g., CNNs and ViTs. As a student network, the ViT-based networks leverages its global attention mechanism to attain the comparable performance with more realistic images. However, as a teacher network, ViT-based networks does not have strong ability to teach the student networks, yielding weak performance. Additionally, due to the low-pass filter characteristics of ViTs, effective learning becomes challenging when the sample size is limited and class-specific features are incomplete.
4.4 Training Efficiency